Thomas’ Calculus Early Transcendentals 13ed

Study guide for quiz 05

Quiz problems include both the lecture contents and homework problems.

1. Classify possible locations of local or global minimum or maximum for a continuous function on a closed interval.

2. Study the statements, proofs and applications of Rolle’s Theorem and Mean Value Theorem. Be careful about the difference between Intermediate Value Theorem and Mean Value Theorem.

3. Study how to determine whether a critical point is a local minimum, local maximum, or neither using the first derivative test.